

********************************************************************
UARS WEEKLY STATUS REPORT - 27 FEB 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off on Feb. 20 and turned back on Feb. 22. On all
other times during this past week, SUSIM successfully gathered
its daily mid (1.1 nm) resolution, Mg II, occultation, solar
pointing offset and high resolution continuous monitoring scans for
periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003229 has been
reprocessed and archived at the GSFC DAAC. HEPSA data files between
2003230 and 2004269 are waiting for DAAC collection.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1994029. MEPS data files between
1994030 and 1994059 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 14 February 2004 (Orbit 67992, GMT Day 045) through
20 February 2004 (Orbit 68096, GMT Day 051).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 045/1011 67998 FULLTIME TO LOW POWER
PEM 047/2205 68035 POWER OFF
HALOE 048/1003 68043 POWER OFF (EARLY)
SUSIM 051/0018 68082 POWER OFF
SOLSTICE 051/0110 68082 FULLTIME TO DAY-ONLY OPS
The beta angle ranged between the angles of 25.8 degrees to 02.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 21 February 2004 (DOY 052).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.28 degrees C. Additional
battery performance data for each battery on 13 February and 20
February 2004 are provided in Appendix D.
The clock error ranged from 10.9 to -09.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HRDI = 01
PEM = 01
HALOE = 01
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began testing the new version of the CMS code this week. The
testing is scheduled to be completed by 05 March 2004. A tentative
date of 09 March 2004 has been set for a mini-ORR to review the
results of testing and the plans to utilize the new 3-TDRS
functionality.
MEETINGS
The FOT held a meeting on 19 February 2004 to discuss the CMS code
problem and the regression test plan for the following weeks.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 February 2004 (DOY 052).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 20 February 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67998 045/1011 04-054 HRDI FULLTIME TO LOW POWER
68019 046/2057 04-055 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68035 047/2205 04-056 PEM POWER OFF
68038 048/0305 04-057 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68043 048/1003 04-058 HALOE POWER OFF (EARLY DUE TO LOW POWER)
68082 051/0018 04-059 SUSIM POWER OFF
68082 051/0116 04-060 SOLSTICE FULLTIME TO DAY-ONLY OPS
68092 051/1638 04-061 MPS VT LEVEL 6 (YAW)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 20 February 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 20 February 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68019 046/2102 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68038 048/0305 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 February 2004
13 February 2004 - GMT Day 044 Beta = 27.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-089.6 +6.43/+5.14 +28.8/-15.2 24.0 1.04 20.7
3 +11.2/-134.4 +5.14/+3.87 +28.0/-15.2 24.0 1.03 20.5
20 February 2004 - GMT Day 051 Beta = 02.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-44.8 +5.14/+3.87 +28.4/-12.8 25.8 1.04 20.0
3 +11.2/-50.4 +3.87/+2.60 +27.2/-12.8 25.6 1.04 19.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 20 February 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-001 FOT See List Jan 04 30832 Generic Late Acquisition
04-009 FOT 67854 035/2126 N/A EPHEMERIS REJECTED
04-010 FOT 67934 041/0409 35016 4 Sec Data Dropout
04-012 FOT 67965 043/0643 35015 Late Acq & No GCMRs/UPDs
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-011 FOT See List Feb 04 30832 Generic Late Acquisition